what is 72 out of 85 expressed as a percentage please?
Answer by stanbon(75887)

what is 72 out of 85 expressed as a percentage please?
----
Solve for "x":
x/100 = 72/85
---
x = 100(72/85)
x = 84.71%
======================
Cheers,
Stan H.
======================